# Spin of fractional quantum Hall neutral modes and “missing states” on a sphere

### v1

@@ Abstract
A low-energy neutral quasiparticle in a fractional quantum Hall system appears in the latter’s energy spectrum on a sphere as a series of many-body excited states labeled by the angular momentum $L$ and whose energy is a smooth function of $L$ in the limit of large sphere radius. We argue that the signature of a nonvanishing spin (intrinsic angular momentum) $s$ of the quasiparticle is the absence, in this series, of states with total angular momentum less than $s$. We reinterpret the missing of certain states, observed in an exact-diagonalization calculation of the spectrum of the $\nu=7/3$ FQH state in a wide quantum well as well as in many proposed wave functions for the excited states as a consequence of the spin-2 nature of the zero-momentum magnetoroton.
@@ 2 The puzzle of missing states
Consider a gapped FQH state. Let us assume that the system possesses a quasiparticle with a dispersion relation that corresponds to a gap $E_{0}$ at zero momentum and an effective mass $m$,
@@
In the FQH context, only electrically neutral particles can be assigned a dispersion relation. (Equation (2.1), thus, cannot be applied to, e.g., the Laughlin quasihole or quasielectron.) For simplicity, let us assume that the quasiparticle (2.1) has a much lower energy than any other excitations in the system, although much of the discussion that follows is valid under a much weaker condition.
@@
For concreteness, we focus on one particular quantum Hall system, namely, the $\nu=1/3$ state. The lowest-energy neutral quasiparticle in this case is the magnetoroton [3]. For the $\nu=1/3$ FQH state with Coulomb interaction the magnetoroton dispersion relation has a minimum at $k_{0}\sim 1/\ell_{B}$ (hence the name) and a local maximum at $\mathbf{k}=0$. The magnetoroton at $\mathbf{k}=0$ is, however, buried in a continuum of states, which presumably can be understood as two-quasiparticle states containing two magnetorotons near the minimum of the dispersion curve. The long-wavelength ($\mathbf{k}\approx 0$) magnetoroton, thus, should be unstable against decay into two finite-momentum magnetorotons. The instability of the long-wavelength magnetoroton is, however, not an unavoidable consequence of the FQH effect. In fact, numerical simulation of the $\nu=7/3$ state (i.e., $\nu=1/3$ state on the second Landau level) in a wide quantum well [12] reveal a regime where the magnetoroton at $\mathbf{k}=0$ is stable, with an energy less than twice the energy at the minimum at $\mathbf{k}\neq 0$, and with a positive effective mass. One expects that in this case the series of states (2.2) and (2.3) can be reproduced numerically.
@@ 3 The main argument
The Berry phase makes the problem equivalent to that of a particle moving on a sphere under the influence of the magnetic field of a magnetic monopole located at the center of the sphere. We emphasize that this magnetic monopole is fictitious and unrelated to the source of the magnetic field that gives rise to the FQHE. The quasiparticle under consideration is charge-neutral with respect to the physical electromagnetic field, but its spin makes it “feels” a fictitious magnetic field.
@@
[…]
@@
States with $L<|s|$ do not appear in the spectrum. Thus, the absence of $L=0$ and $L=1$ modes (and the presence of $L\geq 2$ modes) observed in Ref. [12] is not a finite-size effect, but a direct evidence showing that spin of the long-wavelength magnetoroton is 2 (more precisely, $+2$ or $-2$).
@@ 4 Conclusion
We note that before Ref. [12], the lack of the $L=1$ excited state has already been encountered in many studies. In particular, in the composite-fermion framework, the magnetoroton is obtained by moving a composite fermion from a lower, occupied $\Lambda$ level to a higher, unoccupied one, and then projecting the resulting wavefunction to the lowest Landau level. In the $\nu=1/3$ case, one can create a state with $L=1$ by promoting a composite fermion from the lowest $\Lambda$-level to the next $\Lambda$-level, but that state annihilated, in a nontrivial manner, by the projection to the lowest Landau level[13, 14]. The absence of the $L=1$ mode is also a feature of analytic constructions of the excited states directly on the lowest Landau level [15, 16]
@@
It has been argued that Jain states with $\nu=n/(2n\pm 1)$ accommodate not only spin-2 but also spin-3, spin-4, etc., excitations [22, 23]. If a spin-$s$ mode is a well-defined excitation, one should observe the missing of corresponding modes with $L<s$ on the sphere. This may be related to the “exclusion rules” for counting excited states in the composite fermion picture [14].
@@ A.1 Metric and notation
Note that since we are working on a static curved space in $2+1$D, the spin connection defined in equations (A.4) and (A.5) correspond to $\omega_{i}^{12}=-\omega_{i}^{21}$ in the usual notation of the general geometry [27, 28]. The spin connection (A.5) also satisfies the relation to the Ricci curvature of a sphere with radius $R$
@@
From the above convention, on a static curved space, the covariant derivative acting on the neutral field $\Psi$ with the $-s$ representation of rotational group $SU(2)$ is given by
@@
$\Sigma_{ab}$ is the component of spin operator in spin-$s$ representation in the direction perpendicular to the plane $(a,b)$. Since we have the only non-zero component of spin connection in the static curved space in $2+1$D are $\omega_{i}^{12}=-\omega_{i}^{21}=\omega_{i}$, only $\Sigma_{12}=-\Sigma_{21}$ appear in the covariant derivative (A.7).
@@ A.2 Energy spectrum of a spin-$s$ excitation on a sphere
[…]
@@
which behaves as the usual angular momentum operator with the commutation relation $[J_{i},J_{j}]=i\epsilon^{ijk}J_{k}$. We can check explicitly that the Hamiltonian is rotational invariant $[J_{i},H]=0$. Therefore eigenstates of the Hamiltonian $H$ can be chosen to have well-defined total angular momentum,
@@
which is Eq. (3.2) in the main text. Furtheremore, since the Hamiltonian is rotational invariant, the degeneracy of each energy level is $12L+$ 1. Since the momentum operator $\vec{\Pi}=-i\vec{\nabla}_{s}-e\vec{A}$ is Hermitian, it is obviously that
$\langle\Psi_{L}|\left(-i\vec{\nabla}_{s}-e\vec{A}\right)^{2}|\Psi_{L}\rangle\geq 0,$ (A.19)

### v2

# Spin of fractional quantum Hall neutral modes and ``missing states'' on a sphere

@@ Abstract
A low-energy neutral quasiparticle in a fractional quantum Hall system appears in the latter's energy spectrum on a sphere as a series of many-body excited states labeled by the angular momentum $L$ and whose energy is a smooth function of $L$ in the limit of large sphere radius. We argue that the signature of a nonvanishing spin (intrinsic angular momentum) $s$ of the quasiparticle is the absence, in this series, of states with total angular momentum less than $s$. We reinterpret the missing of certain states, observed in an exact-diagonalization calculation of the spectrum of the $\nu=7/3$ FQH state in a wide quantum well as well as in many proposed wave functions for the excited states as a consequence of the spin-2 nature of the zero-momentum magnetoroton.
@@ 2 The puzzle of missing states
Consider a gapped FQH state. Let us assume that the system possesses a quasiparticle with a dispersion relation that corresponds to a gap $E_{0}$ at zero momentum and an effective mass $m$  [^2] The dispersion relation in Eq. (2.1) is chosen to respect rotational symmetry. As demonstrated in the Appendix, the specific form of the dispersion does not affect the main conclusion regarding the missing state, as long as rotational symmetry is maintained.,
@@
In the context of the fractional quantum Hall effect, only electrically neutral excitations can be assigned a dispersion relation, since, in two-dimensional systems in a magnetic field, charged excitations do not have well-defined momentum (Equation (2.1), thus, cannot be applied to, e.g., the Laughlin quasihole or quasielectron.) For simplicity, let us assume that the quasiparticle (2.1) has a much lower energy than any other excitations in the system, although much of the discussion that follows is valid under a much weaker condition.
@@
For concreteness, we focus on one particular quantum Hall system, namely, the $\nu=1/3$ state. The lowest-energy neutral quasiparticle in this case is the magnetoroton [3]. For the $\nu=1/3$ FQH state with Coulomb interaction the magnetoroton dispersion relation has a minimum at $k_{0}\sim 1/\ell_{B}$ (hence the name) and a local maximum at $\mathbf{k}=0$. The magnetoroton at $\mathbf{k}=0$ is, however, buried in a continuum of states, which presumably can be understood as two-quasiparticle states containing two magnetorotons near the minimum of the dispersion curve. The long-wavelength ($\mathbf{k}\approx 0$) magnetoroton, thus, could be unstable against decay into two finite-momentum magnetorotons. The instability of the long-wavelength magnetoroton is, however, not an unavoidable consequence of the FQH effect. In fact, numerical simulation of the $\nu=7/3$ state (i.e., $\nu=1/3$ state on the second Landau level) in a wide quantum well [12] reveal a regime where the magnetoroton at $\mathbf{k}=0$ is stable, with an energy less than twice the energy at the minimum at $\mathbf{k}\neq 0$, and with a positive effective mass. One expects that in this case the series of states (2.2) and (2.3) can be reproduced numerically.
@@ 3 The main argument
The Berry phase makes the problem equivalent to that of a particle moving on a sphere under the influence of the magnetic field of a magnetic monopole located at the center of the sphere. We emphasize that this magnetic monopole is fictitious and unrelated to the source of the magnetic field that gives rise to the FQHE. The quasiparticle under consideration is charge-neutral with respect to the physical electromagnetic field, but its spin makes it ``feels'' a fictitious magnetic field.
@@
where $L$ is the angular momentum. It can be demonstrated that $L(L+1)-(eg)^{2}\geq 0$ [^4] For the detailed derivation, see the Appendix.. Consequently, unlike the situation without a magnetic monopole, here $L$ cannot take arbitrary integer values, but only
@@
States with $L<|s|$ do not appear in the spectrum. Thus, the absence of $L=0$ and $L=1$ modes (and the presence of $L\geq 2$ modes) observed in Ref. [12] is not a finite-size effect, but a direct evidence showing that spin of the long-wavelength magnetoroton is 2 (more precisely, +2+2 or 2-2).
@@ 4 Conclusion
We note that before Ref. [12], the lack of the $L=1$ excited state has already been encountered in many studies. In particular, in the composite-fermion framework, the magnetoroton is obtained by moving a composite fermion from a lower, occupied $\Lambda$ level to a higher, unoccupied one, and then projecting the resulting wavefunction to the lowest Landau level. In the $\nu=1/3$ case, one can create a state with $L=1$ by promoting a composite fermion from the lowest $\Lambda$-level to the next $\Lambda$-level, but that state is annihilated, in a nontrivial manner, by the projection to the lowest Landau level[13, 14]. The absence of the $L=1$ mode is also a feature of analytic constructions of the excited states directly on the lowest Landau level [15, 16]
@@
It has been argued that Jain states with $\nu=n/(2n\pm 1)$ accommodate not only spin-2 but also spin-3, spin-4, etc., excitations  [22, 23]. If a spin-$s$ mode is a well-defined excitation, one should observe the missing of corresponding modes with $L<s$ on the sphere. This may be related to the ``exclusion rules'' for counting excited states in the composite fermion picture [14]. Using the exclusion rules proposed in [14], one finds exactly 2(L1)2(L-1) states with angular momentum for $L\geq 2$, constructed from composite fermions. This state counting is consistent with the existence of two neutral modes for each spin $s\geq 2$. However, it remains unclear whether the exclusion rules in [14] are complete. Therefore, the complete count of neutral modes for each spin at filling fraction $\nu=2/5$ is still an open question.
For the Jain states at filling $\nu=n/(2n\pm 1)$, in the large-$n$ limit, we expect a single spin-$s$ excitation for each $s\geq 2$, corresponding to deformations of the composite Fermi surface [23]. In contrast, for the Jain states at $\nu=\frac{n}{4n\pm 1}$, we anticipate the presence of an additional spin-2 mode [7, 8, 9] in the large $n$ limit.
@@ A.1 Metric and notation
Note that since we are working on a static curved space in 2+12+1D, the spin connection defined in equations (A.4) and (A.5) correspond to $\omega_{i}^{12}=-\omega_{i}^{21}$ in the usual notation of the general geometry [27, 28]. The spin connection (A.5) also satisfies the relation to the Ricci curvature of a sphere with radius $R$
@@
From the above convention, on a static curved space, the covariant derivative acting on the neutral field $\Psi$ with the spins-s representation of rotational group $SU(2)$ is given by
@@
$\Sigma_{ab}$ is the component of spin operator in spin-$s$ representation in the direction perpendicular to the plane $(a,b)$. Since we have the only non-zero component of spin connection in the static curved space in 2+12+1D are $\omega_{i}^{12}=-\omega_{i}^{21}=\omega_{i}$, only $\Sigma_{12}=-\Sigma_{21}$ appear in the covariant derivative (A.7).
@@ A.2 Energy spectrum of a spin-$s$ excitation on a sphere
$\mathcal{S}=\int dtd\mathbf{\theta}d\phi\sqrt{g}\left\{i\Psi^{\dagger}\partial_{t}\Psi-\Delta\Psi^{\dagger}\Psi-\frac{g^{ij}}{2m}\left(\partial_{i}-is\omega_{i}\right)\Psi^{\dagger}\left(\partial_{j}+is\omega_{j}\right)\Psi\right\},$ (A.9)
@@
which behaves as the usual angular momentum operator with the commutation relation $J_{i},J_{j}]=i\epsilon^{ijk}J_{k}$. We can check explicitly that the Hamiltonian is rotational invariant $J_{i},H]=0$. Therefore eigenstates of the Hamiltonian $H$ can be chosen to have well-defined total angular momentum,
@@
which is Eq. (3.2) in the main text. Furtheremore, since the Hamiltonian is rotational invariant, the degeneracy of each energy level is $2L+1$. Since the momentum operator $\vec{\Pi}=-i\vec{\nabla}_{s}-e\vec{A}$ is Hermitian, it is obviously that
$\langle\Psi_{L}|(\vec{\Pi})^{2}|\Psi_{L}\rangle=\langle\Psi_{L}|\left(-i\vec{\nabla}_{s}-e\vec{A}\right)^{2}|\Psi_{L}\rangle\geq 0,$ (A.19)
@@
Some comments are in order, given that the momentum operator $\vec{\Pi}$ is Hermitian, the constraint (A.19) is exact, given that the excitation has a well-defined angular momentum. Therefore, the main conclusion (3.3) does not depend on the explicit dispersion relation of the gapped neutral excitation, given that the dispersion, and therefore the effective Hamiltonian, is rotational invariant.